OpenStudy (anonymous):

After 4 years, Aspen earned $510 in simple interest from a CD into which she initially deposited $3000. What was the annual interest rate of the CD? A. 13% B. 8.5% C. 4.25% D. 11.75% I=_P_∗_R∗_T_ 100 sorry about the formula i know it look crappy :P

OpenStudy (anonymous):

@kl0723

OpenStudy (anonymous):

do u know how to do it?

OpenStudy (kl0723):

is that the following? \[I = \frac{ P \times R \times T }{ 100 }\]

OpenStudy (anonymous):

yep

OpenStudy (kl0723):

ok so (I) is the interest money that she earned, then (P) is the principal amount, (R) is the rate in decimal form, (T) is the time the money was invested, so from this equation we need to solve for R :)

OpenStudy (anonymous):

i know that P=$3000, I=$510, T=4 n R=? i have everything written i just need to know how to solve it

OpenStudy (kl0723):

remove the 100 from the bottom, you do not need it

OpenStudy (kl0723):

the original formula is I = PRT

OpenStudy (kl0723):

so to solve for R we need to divide the right side by P and T, doing the same on the left side can you draw that step?

OpenStudy (anonymous):

that weird cuz they told me it also involves the 100 but what lets just go with urs

OpenStudy (kl0723):

the 100 is to get a direct answer, we will find R in decimal form and then convert it to percent form so you know how to that as well :)

OpenStudy (anonymous):

yea i don't know what u mean

OpenStudy (kl0723):

lets isolate R to get our answer, can you do that?

OpenStudy (kl0723):



OpenStudy (kl0723):

since P,R and T are factors, that means they are multiplying, we need to divide in order to leave one of them alone

OpenStudy (anonymous):

my apologize i don't know how to do these types of question cuz im only 13 n i have to do these high school work

OpenStudy (kl0723):

np that's what you're here for :)

OpenStudy (kl0723):